A 2.69 kg sphere makes a perfectly inelastic collision with a second sphere that is initially at rest. The composite system moves with a speed equal to one fifth the original speed of the 2.69 kg sphere. What is the mass of the second sphere?   

Plese help!!
Answer #1

m_{1}V_{o1}+m_{2}V_{o2}=(m_{1}+m_{2})V

m_{1}V_{o1}=(m_{1}+m_{2})V

Isolating the mass m_{2},

m_{2}=m_{1}\left ( \frac{V_{o1}}{V}-1 \right )=m_{1}\left ( \frac{5V}{V}-1 \right )=4m_{1}=4(2.69Kg)

or

{\color{Blue} m_{2}=10.8\, Kg}